OpenStudy (anonymous):

find the measures of an angle and its complement if one angle measures 24 degrees more than the other.

OpenStudy (anonymous):

complements sums to 90. 90=24+x+x =24+2x 66=2x >>>>>x=33 >>>>>24+x=57